solving for "y"
x+2y=6
x-3y=12
2x-5y=0
2x-3y=-3
4x+y=3
3x-2y=-8

Answer by waynest(281)

1.) x + 2y = 6
x - x + 2y = -x + 6
2y = -x + 6
2y/2 = -x/2 + 6/2
y = -x/2 + 3
2.) x - 3y = 12
x - 3y + 3y -12 = 12 - 12 + 3y
x - 12 = 3y
x/3 - 12/3 = 3y/3
x/3 - 4 = y
3.) 2x - 5y = 0
2x - 5y + 5y = 5y
2x/5 = 5y/5
2x/5 = y
4.) 2x - 3y -3
2x - 3y + 3y + 3 = - 3 + 3 + 3y
2x + 3 = 3y
2x/3 + 3/3 = 3y/3
2x/3 + 1 = y
5.) 4x + y = 3
4x - 4x + y = 3 - 4x
y = 3 - 4x
6.) 3x - 2y = -8
3x - 2y + 2y + 8 = -8 + 8 + 2y
3x + 8 = 2y
3x/2 + 8/2 = 2y/2
3x/2 + 4 = y